OpenStudy (kdhdjhio):

Convert the radian measure to degree measure. Use the value of π found on a calculator, and round answers to two decimal places. negative nineteen divided by six, pi

OpenStudy (photon336):

all you have to do is just multiply b the conversion factor: \[\frac{ 180^{o} }{ \pi }*(\pi~radians) = Degrees \]